OpenStudy (anonymous):

I need step by step help for answer asap (2,b) and (4,6) is parallel to the line joining (1,4) and (-1,7). find b.

OpenStudy (anonymous):

First Find the slope

OpenStudy (anonymous):

to the line that is parallel to the line (4,6) and (2,8)

OpenStudy (anonymous):

i get 1/-3 for that

OpenStudy (anonymous):

then for b. i get 6-b/2

OpenStudy (anonymous):

how do i combine to get b?

OpenStudy (anonymous):

Hmm.. well for starters the slope isn't 1/-3 for (-1,7) and (1,4)

OpenStudy (anonymous):

do you know the slope formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well the slope formula is (\[\Delta Y/ \Delta X\]

OpenStudy (anonymous):

so Y2 - Y1 / X2 - X1

OpenStudy (anonymous):

in this case ...

OpenStudy (anonymous):

in (1,4) 1 is the X1 and 4 is the Y1

OpenStudy (anonymous):

and in(-1,7) -1 is the X2 and 7 is Y2

OpenStudy (anonymous):

3/-2 for sloppe

OpenStudy (anonymous):

slope*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we know that the line foe (2,b) AND (4,6) is parallel to (1,4) and (-1,7)

OpenStudy (anonymous):

So that mean that they share the same slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So far in our equation we have Y = 3/-2 +B

OpenStudy (anonymous):

so would i get 3/-2+3/2 = 9 so b=9

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

\[Y= 3/-2X +b\] If we take (4,6) and plug the 4 in the x's place and the 6 in the y's place what would b need to be in order to make that happen

OpenStudy (anonymous):

*\[y=(3/-2)x+b\]

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

\[y=(3/-2)x\] replace the y with 6 and x with 4 \[(6)=(3/-2)(4)+b\]

OpenStudy (anonymous):

y=-6x+b

OpenStudy (anonymous):

6=-6x+b

OpenStudy (anonymous):

No not exactly

OpenStudy (anonymous):

y=(3/−2)x+b replace the y with 6 and x with 4 (6)=(3/−2)(4)+b

OpenStudy (anonymous):

you should get

OpenStudy (anonymous):

6=6+b

OpenStudy (anonymous):

then subtract 6 from both sides?

OpenStudy (anonymous):

*6=-6+b

OpenStudy (anonymous):

Then add sorry i made a typo D:

OpenStudy (anonymous):

12

OpenStudy (anonymous):

SO what we have so far is 6=-6+b

OpenStudy (anonymous):

yes b will be 12

OpenStudy (anonymous):

but we arent done yet

OpenStudy (anonymous):

so our equation should be

OpenStudy (anonymous):

y=(3/-2)x+12

OpenStudy (anonymous):

now we need to figure out b for (2,b) so we place the x with 2 and replace with y with b and see the out come for b

OpenStudy (anonymous):

so what we have now is

OpenStudy (anonymous):

b=(3/-2)(2)+12

OpenStudy (anonymous):

b=9 :)

OpenStudy (anonymous):

Yesh XD

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:3 Good job ;O